# Reformulating Conversational Recommender Systems as Tri-Phase Offline Policy Learning

### v1

@@ Abstract.
Existing Conversational Recommender Systems (CRS) predominantly utilize user simulators for training and evaluating recommendation policies. These simulators often oversimplify the complexity of user interactions by focusing solely on static item attributes, neglecting the rich, evolving preferences that characterize real-world user behavior. This limitation frequently leads to models that perform well in simulated environments but falter in actual deployment. Addressing these challenges, this paper introduces the Tri-Phase Offline Policy Learning-based Conversational Recommender System (TPCRS), which significantly reduces dependency on real-time interactions and mitigates overfitting issues prevalent in traditional approaches. TPCRS integrates a model-based offline learning strategy with a controllable user simulation that dynamically aligns with both personalized and evolving user preferences. Through comprehensive experiments, TPCRS demonstrates enhanced robustness, adaptability, and accuracy in recommendations, outperforming traditional CRS models in diverse user scenarios. This approach not only provides a more realistic evaluation environment but also facilitates a deeper understanding of user behavior dynamics, thereby refining the recommendation process.
@@ 1. Introduction
Conversational Recommender Systems (CRS) have revolutionized traditional recommendation systems by enabling dynamic and interactive conversations with users to elicit fine-grained preferences and provide more accurate and explainable recommendations (Gao et al., 2021). The effectiveness of CRS hinges critically on the availability and quality of real-time user interaction data (Gao et al., 2022; Christakopoulou et al., 2016), which poses significant challenges—data collection is often expensive and intrusive, potentially degrading the user experience. Traditionally, as illustrated in Figure 1(a), CRS models have relied heavily on user simulators during both training and evaluation phases to mimic real-world interactions. However, this reliance on the exact same simulator for both training and evaluation is problematic. It results in the leakage of evaluation rules and information to the policy during the training phase. Ideally, policy learning should be conducted offline, where the policy cannot interact with or explore the environment or users during training (Levine et al., 2020). Instead, it should rely solely on logged data for learning. Furthermore, existing simulators predominantly focus on static user preferences and fail to account for the evolving and individualized nature of user preferences. Consequently, systems trained and evaluated using such simulators perform admirably in controlled, simulated environments but often falter when deployed in real-world settings where user preferences are neither static nor predictable (Gao et al., 2023b).
@@
Recognizing these challenges, we reformulate CRS as a Tri-Phase Offline Policy Learning-based Conversational Recommender System (TPCRS). It is a novel CRS architecture designed to drastically reduce reliance on real-time user interactions and enhance the robustness and accuracy of CRS models. TPCRS integrates a model-based offline policy learning strategy with a controllable user simulation that dynamically adapts to both historical and evolving user preferences. This approach not only addresses the critical flaws of overfitting and unrealistic preference modeling inherent in traditional CRS but also provides a more realistic evaluation environment through its advanced simulation techniques.
@@
As depicted in Figure 1(b), the TPCRS framework consists of three key components: 1) User Model Training, which estimates user preferences from offline data and captures the dynamic and personalized nature of user preferences; 2) Policy Learning, which employs the learned user model as a simulated environment to facilitate the training of recommendation policies; and 3) Policy Evaluation, which assesses the trained policies using an independent, black-box simulator that dynamically adjusts to user interactions for more realistic evaluation. By decoupling the training and evaluation phases, TPCRS aims to mitigate the overfitting issue associated with traditional simulator-based approaches. Additionally, the explicit modeling of user preference dynamics and personalization within the user model can help CRS policies better capture real-world user behavior, leading to improved recommendation performance when deployed in actual user environments.
@@
Building upon the overall TPCRS framework, we now describe the specific implementation details of each key component:
@@
1. Preference Estimation User Model: This component employs a multi-task learning approach to capture and predict user preferences from a rich dataset of offline interactions, comprising a State Tracker for maintaining a comprehensive representation of user preference state, and a Preference Ranking for jointly optimizing item-level and attribute-level preference prediction.
@@
  • A Paradigm Shift in CRS Training and Evaluation: By decoupling the training and evaluation phases, TPCRS introduces a novel offline policy learning framework that is both efficient and scalable, addressing the critical flaws of overfitting and unrealistic preference modeling inherent in traditional CRS approaches.
  • More Realistic Evaluation Environment: The Controllable User Simulation module in TPCRS transcends the static, item-centric assumptions of previous simulators, embracing the personalized and evolving nature of user preferences during interactions. This allows for a comprehensive assessment of policy performance in diverse and realistic user scenarios.
@@ 2. Related Work
Conversational Recommender Systems (CRS) aim to dynamically interact with users to elicit their preferences and provide personalized recommendations (Gao et al., 2021). The research in CRS can be broadly divided into two approaches: policy optimization-based methods (Christakopoulou et al., 2018; Sun and Zhang, 2018; Lei et al., 2020a, b; Deng et al., 2021; Zhang et al., 2022, 2023; Li et al., 2021; Chu et al., 2023; Yang et al., 2024) and dialogue generation-based methods (Li et al., 2018; Chen et al., 2019; Zhou et al., 2020a, b; Li et al., 2022). Policy Optimization-based Methods emphasize real-time, dynamic capturing of user preferences through strategic interactions (Lei et al., 2020b; Liao et al., 2018; Wu et al., 2023). Reinforcement learning techniques (Lei et al., 2020b; Deng et al., 2021) are commonly employed to develop effective recommendation policies. These methods focus on maximizing long-term user satisfaction by learning to ask purposeful questions that reveal user preferences effectively (Christakopoulou et al., 2018; Sun and Zhang, 2018). Dialogue Generation-based Methods use natural language processing (NLP) techniques to generate fluent and contextually appropriate dialogues (Li et al., 2018; Chen et al., 2019; Zhou et al., 2020a; Liao et al., 2022; Dao et al., 2024). While these methods can create natural interactions, they often struggle with decision-making capabilities (Friedman et al., 2023). The reliance on large dialogue corpora limits their ability to handle specific user contexts and detailed preference modeling (Li et al., 2018).
@@
Our TPCRS framework builds upon these offline RL methodologies by introducing a tri-phase policy learning approach for CRS. This approach decouples the training and evaluation phases, leveraging a model-based user simulator to enhance policy robustness and accuracy without the dependency on real-time user data (Yu et al., 2024; Shi et al., 2024).
User Simulators play a critical role in training and evaluating CRS policies by providing synthetic user feedback (Lei et al., 2020a). The design and realism of these simulators significantly impact the effectiveness of the trained policies (Gao et al., 2023b, 2022). Traditional CRS user simulators often assume static user preferences tied directly to item attributes (Lei et al., 2020b; Zhang et al., 2022). These simulators generate feedback based solely on predefined rules, which can lead to overly optimized policies that perform well in simulations but fail to generalize to real-world interactions. Recent advancements advocate for more realistic simulators (Zhang et al., 2023; Zhang, 2023; Deng et al., 2024) that capture the dynamic and personalized nature of user preferences. These simulators model user behavior as evolving and context-dependent, providing a more comprehensive evaluation environment (Zhang and Balog, 2020) for CRS policies. Our TPCRS framework introduces a controllable user simulation module that dynamically aligns with both historical personal preferences and target items preferences.
@@ 3. Preliminary
Figure 3. The Tri-Phase Offline Policy Learning-based Conversational Recommender System (TPCRS) Framework. Preference Estimation User Model (PEUM) is first trained on offline data to capture dynamic and personalized user preferences. The learned user model then serves as the simulated environment for Policy Learning, where the recommendation policy is optimized to maximize long-term user satisfaction. Finally, the trained policy is evaluated using an independent, controllable user simulator that can adapt to diverse user preference scenarios, enabling a comprehensive assessment of the policy’s adaptability.
@@ 4. METHODOLOGY
4.1. Preference Estimation User Model
The Preference Estimation User Model adopts a multi-task learning approach, consisting of a State Tracker module and a Preference Ranking module. The key objective is to develop a comprehensive representation of user preference state that can be leveraged by the policy learning process.
@@ 4.1.1. Data Generation for User Model
To facilitate robust training of the User Model within the TPCRS framework, the Data Generation module constructs a supervised dataset $(X,Y)$ using offline user-item pairs $(u,v)$. This dataset is crucial for effectively simulating and learning user preferences based on historical interactions. The input feature $X$ incorporates user $u$ along with attributes $P_{\text{click}}$ and $P_{\text{nonclick}}$, which are sampled randomly from the attributes of all items previously interacted with by the user. This sampling approach promotes diversity in training data.
@@ 4.1.3. Preference Ranking
By integrating the State Tracker and Preference Ranking modules, the Preference Estimation User Model is designed to capture the complexity of user preferences, enabling it to serve as a robust training environment for the subsequent Conversational Policy Learning component of the TPCRS framework.
@@ 4.2.1. State Representation
The state $\mathbf{s}_{t}$ at interaction $t$ in our reinforcement learning (RL) framework is derived from the State Tracker of the Preference Estimation User Model in Equation (5), capturing the current evolving user preferences. The output of the User Model’s MLP encoder serves as the perceived state of the user’s preference, providing a rich representation for decision-making.
@@ 4.2.2. Action Space
Given the potentially vast action space in CRS, efficient exploration strategies are crucial (Deng et al., 2021). We employ the Preference Ranking of the Preference Estimation User Model in Equation (6) and Equation (7) to prune the candidate item and attribute space, leveraging learned user preference predictions to narrow down choices. The action space is thus defined by selecting the top-ranked items and attributes based on Preference Ranking’s probability distributions:
@@ 4.2.4. PPO Optimization
    where $r_{t}(\theta)$ is the probability ratio of the new and old policies, and $\hat{A}_{t}$ is the advantage estimate at time $t$, reflecting the relative benefit of the selected action (Schulman et al., 2015).
@@ 4.3. Controllable User Simulation
The TPCRS framework introduces the Controllable User Simulation module, which aims to provide a more realistic and comprehensive evaluation environment for CRS policies.
@@ 5. Experiments
To verify the effectiveness of the Tri-Phase Offline Policy Learning-based Conversational Recommender System (TPCRS) framework, as well as to evaluate the experimental effects under different user-centric preference scenario settings, we use the following research questions to guide our experiments:
  • RQ1. How does the Tri-Phase Offline Policy Learning-based Conversational Recommender System (TPCRS) framework’s performance compare to start-of-the-art baselines?
  • RQ2. What is the impact of key component of the TPCRS framework on the overall performance?
  • RQ3. How does the adaptability of the TPCRS framework to the various user simulation scenarios evaluation?
  • RQ4. What are the trends in user preference evolution during interactions with the TPCRS framework?
@@ 5.2.1. User Simulator for Policy Evaluation
The user simulator is initialized with a tuple $(u,V_{\text{history}},V_{\text{target}})$ , where $u$ denotes a user, $V_{\text{history}}$ represents the set of items that the user has interacted with in the training data, and $V_{\text{target}}$ comprises the target items in the test data. Typically, $V_{\text{target}}$ may contain multiple items, we focus on a single target item $v$ from the $(u,v)$ pair. This selection is intended to simplify the testing framework and provide clarity on the interaction dynamics under study.
@@
Table 2. Statistically significant performance comparison of TPCRS and baselines on Success Rate and Average Turn Metrics
@@
AVPPL 0.412 0.540 0.620 8.83 0.194 0.240 0.390 12.29 0.274 0.370 0.426 10.87
Abs Greedy (PEUM) 0.462 0.546 0.660 8.14 0.076 0.144 0.200 13.41 0.301 0.353 0.381 10.47
Max Entropy (PEUM) 0.357 0.446 0.522 9.46 0.052 0.176 0.293 13.15 0.236 0.297 0.397 11.22
TPCRS 0.431 0.590 0.670 7.58 0.105 0.326 0.415 11.62 0.313 0.381 0.429 10.26
@@ 5.2.2. Baselines
  • AVPPL (Zhang et al., 2023): It considers a modified scenario of the MCR scenario, where the user has vague preference. It dynamically models user click and non-click behavior preferences in each round, viewing the user decision process as uncertain preference tendencies rather than absolute binary decisions (accept or reject).
Group 2: Non-Simulator Methods. This group includes methods that operate without a user simulator, utilizing the TPCRS’s Preference Estimation User Model (PEUM) as recommendation model:
@@
Evaluation Consistency. To ensure a fair and comprehensive evaluation, all baseline methods, including those that require user simulators, are assessed using the same user-centric preference simulator employed in the TPCRS framework. Additionally, key policy parameters, such as the number of attribute queries or item recommendations per turn and the maximum number of interaction rounds, are kept consistent across all methods.
@@ 5.2.3. Training Details
The Preference Estimation User Model was pre-trained using offline training data. The Transformer architecture had 4 layers, with a negative sampling size of 2000 for item prediction. The batch size was set to 512, with a hidden size and embedding dimensions of 64 for users, items, and attributes. The learning rate was 0.001, using the Adam optimizer. In the multi-task training, the item prediction loss weight $\lambda$ was set to 0.8.

### v2

@@ Abstract.
Existing Conversational Recommender Systems (CRS) predominantly utilize user simulators for training and evaluating recommendation policies. These simulators often oversimplify the complexity of user interactions by focusing solely on static item attributes, neglecting the rich, evolving preferences that characterize real-world user behavior. This limitation frequently leads to models that perform well in simulated environments but falter in actual deployment. Addressing these challenges, this paper introduces the Tri-Phase Offline Policy Learning-based Conversational Recommender System (TCRS), which significantly reduces dependency on real-time interactions and mitigates overfitting issues prevalent in traditional approaches. TCRS integrates a model-based offline learning strategy with a controllable user simulation that dynamically aligns with both personalized and evolving user preferences. Through comprehensive experiments, TCRS demonstrates enhanced robustness, adaptability, and accuracy in recommendations, outperforming traditional CRS models in diverse user scenarios. This approach not only provides a more realistic evaluation environment but also facilitates a deeper understanding of user behavior dynamics, thereby refining the recommendation process.
@@ 1. Introduction
Conversational Recommender Systems (CRS) have revolutionized traditional recommendation systems by enabling dynamic and interactive conversations with users to elicit fine-grained preferences and provide more accurate and explainable recommendations (Gao et al., 2021). The effectiveness of CRS hinges critically on the availability and quality of real-time user interaction data (Gao et al., 2022; Christakopoulou et al., 2016), which poses significant challenges—data collection is often expensive and intrusive, potentially degrading the user experience. Traditionally, as illustrated in Figure 1(a), CRS models have relied heavily on user simulators during both training and evaluation phases to mimic real-world interactions. However, this reliance on the exact same simulator for both training and evaluation is problematic. It results in the leakage of evaluation rules and information to the policy during the training phase. Ideally, policy learning should be conducted offline, where the policy cannot interact with or explore the environment or users during training (Gao et al., 2023a). Instead, it should rely solely on logged data for learning. Furthermore, existing simulators predominantly focus on static user preferences and fail to account for the evolving and individualized nature of user preferences. Consequently, systems trained and evaluated using such simulators perform admirably in controlled, simulated environments but often falter when deployed in real-world settings where user preferences are neither static nor predictable (Gao et al., 2023b).
@@
Recognizing these challenges, we reformulate CRS as a Tri-Phase Offline Policy Learning-based Conversational Recommender System (TCRS). It is a novel CRS architecture designed to drastically reduce reliance on real-time user interactions and enhance the robustness and accuracy of CRS models. TCRS integrates a model-based offline policy learning strategy with a controllable user simulation that dynamically adapts to both historical and evolving user preferences. This approach not only addresses the critical flaws of overfitting and unrealistic preference modeling inherent in traditional CRS but also provides a more realistic evaluation environment through its advanced simulation techniques.
@@
As depicted in Figure 1(b), the TCRS framework consists of three key components: 1) User Model Training, which estimates user preferences from offline data and captures the dynamic and personalized nature of user preferences; 2) Policy Learning, which employs the learned user model as a simulated environment to facilitate the training of recommendation policies; and 3) Policy Evaluation, which assesses the trained policies using an independent, black-box simulator that dynamically adjusts to user interactions for more realistic evaluation. By decoupling the training and evaluation phases, TCRS aims to mitigate the overfitting issue associated with traditional simulator-based approaches. Additionally, the explicit modeling of user preference dynamics and personalization within the user model can help CRS policies better capture real-world user behavior, leading to improved recommendation performance when deployed in actual user environments.
@@
Building upon the overall TCRS framework, we now describe the specific implementation details of each key component:
@@
1. Conversational User Model: This component employs a multi-task learning approach to capture and predict user preferences from a rich dataset of offline interactions, comprising a State Tracker for maintaining a comprehensive representation of user preference state, and a Preference Ranking for jointly optimizing item-level and attribute-level preference prediction.
@@
  • A Paradigm Shift in CRS Training and Evaluation: By decoupling the training and evaluation phases, TCRS introduces a novel offline policy learning framework that is both efficient and scalable, addressing the critical flaws of overfitting and unrealistic preference modeling inherent in traditional CRS approaches.
  • More Realistic Evaluation Environment: The Controllable User Simulation module in TCRS transcends the static, item-centric assumptions of previous simulators, embracing the personalized and evolving nature of user preferences during interactions. This allows for a comprehensive assessment of policy performance in diverse and realistic user scenarios.
@@ 2. Related Work
Conversational Recommender Systems (CRS) aim to dynamically interact with users to elicit their preferences and provide personalized recommendations (Gao et al., 2021). The research in CRS can be broadly divided into two approaches: policy optimization-based methods (Christakopoulou et al., 2018; Sun and Zhang, 2018; Lei et al., 2020a, b; Deng et al., 2021; Zhang et al., 2022, 2023; Li et al., 2021; Chu et al., 2023; Yang et al., 2024) and dialogue generation-based methods (Li et al., 2018; Chen et al., 2019; Zhou et al., 2020; Liao et al., 2022). Policy Optimization-based Methods emphasize real-time, dynamic capturing of user preferences through strategic interactions (Lei et al., 2020b; Wu et al., 2023). Reinforcement learning techniques (Lei et al., 2020b; Deng et al., 2021) are commonly employed to develop effective recommendation policies. These methods focus on maximizing long-term user satisfaction by learning to ask purposeful questions that reveal user preferences effectively (Christakopoulou et al., 2018; Sun and Zhang, 2018). Dialogue Generation-based Methods use natural language processing (NLP) techniques to generate fluent and contextually appropriate dialogues (Li et al., 2018; Chen et al., 2019; Zhou et al., 2020; Dao et al., 2024). While these methods can create natural interactions, they often struggle with decision-making capabilities (Friedman et al., 2023). The reliance on large dialogue corpora limits their ability to handle specific user contexts and detailed preference modeling (Li et al., 2018).
@@
Our TCRS framework builds upon these offline RL methodologies by introducing a tri-phase policy learning approach for CRS. This approach decouples the training and evaluation phases, leveraging a model-based user simulator to enhance policy robustness and accuracy without the dependency on real-time user data (Yu et al., 2024; Shi et al., 2024).
User Simulators play a critical role in training and evaluating CRS policies by providing synthetic user feedback (Lei et al., 2020a). The design and realism of these simulators significantly impact the effectiveness of the trained policies (Gao et al., 2023b, 2022). Traditional CRS user simulators often assume static user preferences tied directly to item attributes (Lei et al., 2020b; Zhang et al., 2022). These simulators generate feedback based solely on predefined rules, which can lead to overly optimized policies that perform well in simulations but fail to generalize to real-world interactions. Recent advancements advocate for more realistic simulators (Zhang et al., 2023; Zhang, 2023; Deng et al., 2024) that capture the dynamic and personalized nature of user preferences. These simulators model user behavior as evolving and context-dependent, providing a more comprehensive evaluation environment (Zhang and Balog, 2020) for CRS policies. Our TCRS framework introduces a controllable user simulation module that dynamically aligns with both historical personal preferences and target items preferences.
@@ 3. Preliminary
Figure 3. The Tri-Phase Offline Policy Learning-based Conversational Recommender System (TCRS) Framework. Conversational User Model (CUM) is first trained on offline data to capture dynamic and personalized user preferences. The learned user model then serves as the simulated environment for Policy Learning, where the recommendation policy is optimized to maximize long-term user satisfaction. Finally, the trained policy is evaluated using an independent, controllable user simulator that can adapt to diverse user preference scenarios, enabling a comprehensive assessment of the policy’s adaptability.
@@ 4. METHODOLOGY
4.1. Conversational User Model
The Conversational User Model adopts a multi-task learning approach, consisting of a State Tracker module and a Preference Ranking module. The key objective is to develop a comprehensive representation of user preference state that can be leveraged by the policy learning process.
@@ 4.1.1. Data Generation for User Model
To facilitate robust training of the User Model within the TCRS framework, the Data Generation module constructs a supervised dataset $(X,Y)$ using offline user-item pairs $(u,v)$. This dataset is crucial for effectively simulating and learning user preferences based on historical interactions. The input feature $X$ incorporates user $u$ along with attributes $P_{\text{click}}$ and $P_{\text{nonclick}}$, which are sampled randomly from the attributes of all items previously interacted with by the user. This sampling approach promotes diversity in training data.
@@ 4.1.3. Preference Ranking
By integrating the State Tracker and Preference Ranking modules, the Conversational User Model is designed to capture the complexity of user preferences, enabling it to serve as a robust training environment for the subsequent Conversational Policy Learning component of the TCRS framework.
@@ 4.2.1. State Representation
The state $\mathbf{s}_{t}$ at interaction $t$ in our reinforcement learning (RL) framework is derived from the State Tracker of the Conversational User Model in Equation (5), capturing the current evolving user preferences. The output of the User Model’s MLP encoder serves as the perceived state of the user’s preference, providing a rich representation for decision-making.
@@ 4.2.2. Action Space
Given the potentially vast action space in CRS, efficient exploration strategies are crucial (Deng et al., 2021). We employ the Preference Ranking of the Conversational User Model in Equation (6) and Equation (7) to prune the candidate item and attribute space, leveraging learned user preference predictions to narrow down choices. The action space is thus defined by selecting the top-ranked items and attributes based on Preference Ranking’s probability distributions:
@@ 4.2.4. PPO Optimization
    where $r_{t}(\theta)$ is the probability ratio of the new and old policies, and $\hat{A}_{t}$ is the advantage estimate at time $t$, reflecting the relative benefit of the selected action.
@@ 4.3. Controllable User Simulation
The TCRS framework introduces the Controllable User Simulation module, which aims to provide a more realistic and comprehensive evaluation environment for CRS policies.
@@ 5. Experiments
To verify the effectiveness of the Tri-Phase Offline Policy Learning-based Conversational Recommender System (TCRS) framework, as well as to evaluate the experimental effects under different user-centric preference scenario settings, we use the following research questions to guide our experiments:
  • RQ1. How does the Tri-Phase Offline Policy Learning-based Conversational Recommender System (TCRS) framework’s performance compare to start-of-the-art baselines?
  • RQ2. What is the impact of key component of the TCRS framework on the overall performance?
  • RQ3. How does the adaptability of the TCRS framework to the various user simulation scenarios evaluation?
  • RQ4. What are the trends in user preference evolution during interactions with the TCRS framework?
@@ 5.2.1. User Simulator for Policy Evaluation
The user simulator is initialized with a tuple $(u,V_{\text{his}},V_{\text{tar}})$ , where $u$ denotes a user, $V_{\text{his}}$ represents the set of items that the user has interacted with in the training data, and $V_{\text{tar}}$ comprises the target items in the test data. Typically, $V_{\text{tar}}$ may contain multiple items, we focus on a single target item $v$ from the $(u,v)$ pair. This selection is intended to simplify the testing framework and provide clarity on the interaction dynamics under study.
@@
Table 2. Statistically significant performance comparison of TCRS and baselines on Success Rate and Average Turn Metrics
@@
VPPL 0.412 0.540 0.620 8.83 0.194 0.240 0.390 12.29 0.274 0.370 0.426 10.87
Abs Greedy (CUM) 0.462 0.546 0.660 8.14 0.076 0.144 0.200 13.41 0.301 0.353 0.381 10.47
Max Entropy (CUM) 0.357 0.446 0.522 9.46 0.052 0.176 0.293 13.15 0.236 0.297 0.397 11.22
TCRS 0.431 0.590 0.670 7.58 0.105 0.326 0.415 11.62 0.313 0.381 0.429 10.26
@@ 5.2.2. Baselines
  • VPPL (Zhang et al., 2023): It considers a modified scenario of the MCR scenario, where the user has vague preference. It dynamically models user click and non-click behavior preferences in each round, viewing the user decision process as uncertain preference tendencies rather than absolute binary decisions (accept or reject).
Group 2: Non-Simulator Methods. This group includes methods that operate without a user simulator, utilizing the TCRS’s Conversational User Model (CUM) as recommendation model:
@@
Evaluation Consistency. To ensure a fair and comprehensive evaluation, all baseline methods, including those that require user simulators, are assessed using the same user-centric preference simulator employed in the TCRS framework. Additionally, key policy parameters, such as the number of attribute queries or item recommendations per turn and the maximum number of interaction rounds, are kept consistent across all methods.
@@ 5.2.3. Training Details
The Conversational User Model was pre-trained using offline training data. The Transformer architecture had 4 layers, with a negative sampling size of 2000 for item prediction. The batch size was set to 512, with a hidden size and embedding dimensions of 64 for users, items, and attributes. The learning rate was 0.001, using the Adam optimizer. In the multi-task training, the item prediction loss weight $\lambda$ was set to 0.8.